A student’s iPhone 6 (mass = 0.129 kg) is tied to a 40 cm string and rotated at a constant speed of 5.0 m/s in a circular path.

a) What is the magnitude of the iPhone’s centripetal acceleration?


b) Find the magnitude of the centripetal force on the iPhone.
1
Expert's answer
2020-04-06T08:42:50-0400

a)


"a=\frac{v^2}{r}=\frac{5^2}{0.4}=62.5\frac{m}{s^2}"

b)


"F=ma=(0.129)(62.5)=8.06\ N"